OpenStudy (anonymous):

How do I find a Particular Integral(PI), for this equation? : \frac{ d^2y }{ dx^2 } + 16y = x^2 + 3

OpenStudy (anonymous):

I already have the Complementary Function(CF) it is : \[y _{c} = Acos4x + Bsin4x\]

OpenStudy (anonymous):

the equation is actually: \[\frac{ d^2y }{ dx^2 } + 16y = x^2 + 3\]

OpenStudy (anonymous):

ok if u have to find paarticular solution then see followung steps

OpenStudy (anonymous):

You just let the constant C become a variable。

OpenStudy (anonymous):

\[Yp=x ^{2} +3/D ^{2}+16 \] \[Yp=(x ^{2} +3)(D ^{2}+16)^{-1} \] \[(D ^{2}+16)^{-1}=D ^{2}-16D ^{4}+..................\]

OpenStudy (anonymous):

the next terms including second term can be neglected because derivative higher than third of x is 0

OpenStudy (anonymous):

k let me try that now

OpenStudy (anonymous):

Yp=\[(x ^{2}+3)*D ^{2}\] \[D ^{2}x ^{2}+D ^{2}3=2+0\] Yp=2

OpenStudy (anonymous):

This is a nonhomogeneous diffence equation.First of all, you sove the homogeneous equation。

OpenStudy (anonymous):

\[\frac{ d^2y }{dx^2}+16y=0\]

OpenStudy (anonymous):

homogeneous eqn has been solved and I hv the CF I posted it on the 1st line in this timeline

OpenStudy (anonymous):

the Auxillary eqn, had complex roots

OpenStudy (anonymous):

Find the eigenequation.

OpenStudy (anonymous):

fazeelayza.He actually use the infnit series.Both way are ok.

OpenStudy (anonymous):

So why not use the real part of the complex number.

OpenStudy (anonymous):

Or he use the Opeator.Ok ,I don't know.

OpenStudy (anonymous):

\[\alpha = 0,\beta = 4\], since k = 4i

OpenStudy (anonymous):

\[ax^2+bx+c\]plug in the equation to get a b c